OpenStudy (anonymous):

Given f(x) = –5x + 1, what is f(–2)? I give Medals!

hartnn (hartnn):

to get f(-2) just plug in x= -2 in f(x) ! what do u get if you put x = -2 in -5x+1 ?

OpenStudy (anonymous):

f(-2)= (-5*-2)+1=11

OpenStudy (anonymous):

F(x)= -5x+1 -----> F(-2)= -5(-2)+1 = 11

OpenStudy (anonymous):

-5(-2)+1 is eleven i presume

hartnn (hartnn):

11 is correct :)

OpenStudy (anonymous):

yes 11

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome ^_^